OpenStudy (anonymous):

I need a little bit of help.... √(x+2)+4=x

OpenStudy (phi):

I would subtract 4 from both sides to get \[ \sqrt{x+2} = x-4 \]

OpenStudy (phi):

do you know how to "square" both sides?

OpenStudy (anonymous):

sqrt(x+2)^2=x-4^2 sqrt(x+4)=x-16?

OpenStudy (phi):

right idea, but , learn this simple rule: squaring a square root makes the square root go away for example \[ \left(\sqrt{x+2}\right)^2 \text{ becomes } x+2\]

OpenStudy (phi):

square means multiply by itself \[ \sqrt{x+2}\cdot \sqrt{x+2} = x+2\]

OpenStudy (phi):

on the right side (x-4) squared means (x-4)(x-4) so you get x+2 = (x-4)(x-4)

OpenStudy (phi):

I would use FOIL to multiply (x-4)(x-4) See http://www.khanacademy.org/math/algebra/polynomials/multiplying_polynomials/v/multiplying-binomials for how to do this